OpenStudy (anonymous):

What is the slope of a line that passes through the point (-2, 3) and is parallel to a line that passes through (3, 7) and (-2, -8)

OpenStudy (debbieg):

Parallel lines have the SAME slope. So find the slope of the line passing through (3,7) and (-2,-8) using the slope formula. That's your slope for the parallel line. (You don't need the point ON the parallel line just to find it's slope.)

OpenStudy (anonymous):

What's the slope formula?

OpenStudy (anonymous):

Is the answer-3

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (3,7) x1 = 3, and y1 = 7 (-2,-8) x2 = -2 and y2 = -8 now just sub that into your formula and that will be your slope

OpenStudy (anonymous):

slope of a parallel line is same . \[slope of line=\frac{ y2-y1 }{x2-x1 }\]

OpenStudy (debbieg):

check your signs. Make sure you are "lining up" the x's and y's of the same point, e.g., whichever y is first in the numerator, you need to use that point's x first in the denominator. If you flip then, you will have the wrong sign in the result.